OpenStudy (anonymous):

Find the function which has the following Laplace transform: 4 / (s^2(s^2 + 4))

OpenStudy (anonymous):

From Mathematica 8 Home Edition: \[\text{InverseLaplaceTransform}\left[\frac{4}{s^2 \left(s^2+4\right)},s,t\right]=4 \left(\frac{t}{4}-\frac{1}{8} \text{Sin}[2 t]\right) \]